# The enhancement of Gaussian mixture modelling as an application to the coherent WaveBurst algorithm in the search for short gravitational wave transients

### v1

@@ I Introduction
Recent work has seen the application of a different ML-based approach to the cWB algorithm Szczepań czyk et al. (2023), in which the authors used the decision tree method of XGBoost to improve the classification of signal and noise transients while maintaining un-modelled requirements to the analysis.
@@ II.3.2 Model optimisation
As mentioned above, a subset of background triggers and simulated WNB triggers are used to construct the background and signal GMM models respectively. While fitting mixture models to each of the two training sets, Gaussian parameters (mean, covariance and weight of each Gaussian) are optimised by the Expectation Maximisation (EM) algorithm Dempster et al. (1977), while the hyper-parameter of number of Gaussian components is specified by optimising detection efficiency with the validation data. During this validation step, models are produced over a range of number of Gaussian components. The log-likelihood values of all triggers (WNB and background) in validation data are calculated for each background and signal model over the range of the number of Gaussian components, allowing for the construction of the $T$ statistic as in eqn. 1. We define detection efficiency as the fraction of WNB validation triggers with $T$ value greater than the detection threshold at a given FAR. In general, the fraction of detected signals will depend on the waveform type and the injected amplitude. But for the detection efficiency calculation here, we use only the WNB triggers and consider all triggers above the detection threshold regardless of the amplitude. For model optimisation, we choose to calculate the detection efficiency at a FAR of 1 in 50 years for each combination of GMM models. The combination of models which give the optimal detection efficiency at the selected FAR threshold are labelled as the GMM models with the optimal number of Gaussian components. An example of the distribution of efficiency at FAR 1 in 50 years over the number of Gaussian components is seen for the O3a analysis in fig 1. Here, GMMs were trained for number of Gaussian components in the range $[140,168]$ for background triggers and in the range $[120,188]$ for the WNB simulations, both varying in steps of 2. The colorbar represents the detection efficiency on validation data for each combination of models. We observe that the detection efficiency varies from $0.48$ to $0.52$ in this range of numbers of Gaussian components, with the maximum occurring at the combination of models with $148$ components for background and $172$ components for WNB simulations.
Previously in Lopez et al. (2022b), the optimal number of Gaussian components had been selected using the Bayesian Information Criterion (BIC). However it was found that this approach did not reliably give the optimal number of Gaussian components to our model.
@@ III.1 Comparisons with old GMM method
  • New method to optimise number of Gaussian components in the models: As demonstrated in Figure 1, this is now done through the optimisation of detection efficiency at an inverse False Alarm Rate (iFAR) threshold of 50 years over combinations of models with differing number of Gaussian components with validation data. Previously, the Bayesian Information Criterion (BIC) was utilized, however was not reliably favouring models with optimal performance.
@@
Figure 2: Percentage increase in $h_{rss50}$ sensitivity for the updated methodology, GMM+, compared to the old GMM methodology seen in Lopez et al. (2022b) for the generic set of ad-hoc waveform injections and CCSN injections at an iFAR threshold of 100 years. Negative increase signifies better sensitivities with the GMM+ methodology.
We directly compare the sensitivity of both methodologies to given waveforms by calculating the percentage increase in $h_{rss50}$ at iFAR $\geq$ 100 years for GMM+ relative to the previous methodology, as seen in Figure 2. A negative percentage increase indicates a lower $h_{rss50}$, and hence an improvement in sensitivity. The results for ad-hoc simulation waveforms are shown in Fig 2(a). Our sensitivities to the majority of Sine-Gaussian waveforms with $Q$=3, $Q$=9 is lowered, however this loss is somewhat expected due to the removal of bias from the models towards these specific injections. Sensitivities to Sine-Gaussian $Q=100$ and White Noise Burst waveforms remain comparable, while GMM+ significantly gains sensitivity to the majority of Gaussian Pulse waveforms. The benefit of eliminating ad-hoc waveform bias from the model is further seen in Figure 2(b), where the consistent ’negative increase’ in $h_{rss50}$ for CCSN injections demonstrates that our new GMM+ methodology increases the data analysis pipeline’s sensitivity to astrophysical source populations.
@@ Acknowledgements.
The authors would like to thank Giovanni Prodi for useful comments. The authors would like to thank the authors of Szczepań czyk et al. (2023) for the production of cWB triggers used for the analysis in this paper. The authors acknowledge the computational resources which aided the completion of this project, provided by LIGO-Laboratory and support by the National Science Foundation (NSF) Grants No..PHY-0757058 and No.PHY-0823459. L.S acknowledges support from the Science and Technology Facilities Council [ST/V506692/1 2446745]. SG acknowledges fellowship support from MHRD, Government of India. J. S. is supported in part by the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2022R1I1A207366012). G.V. acknowledge the support of the National Science Foundation under grant PHY-2207728. I.S.H. was supported by Science and Technology Facilities Council (STFC) grants ST/V001736/1 and ST/V005634/1. AP acknowledges the support from SPARC MoE grant SPARC/2019-2020/P2926/SL, Government of India.
@@ Appendix B CBC events detected by cWB-GMM in O3
Table 4 lists the CBC events detected by GMM+, for both the LH and LHV networks. The triggers are obtained from the entire coincident data used for the O3 cWB all-sky short search.

### v2

# Enhancing search pipelines for short gravitational-wave transients with Gaussian mixture modeling

@@ I Introduction
Recent work has seen the application of a different ML-based approach as post-production to the cWB algorithm Szczepań czyk et al. (2023), in which the authors used the decision tree method, XGBoost, to improve the classification of signal and noise transients while maintaining un-modelled requirements to the analysis. XGBoost learns how to discriminate between typical noise and signal population features through chosen cWB summary statistics, outputting a number between 0 (noise) and 1 (signal) which weights the SNR of given events. As an alternative method to GMM post-production, comparisons of performance are made to the XGBoost approach throughout the paper.
@@ II.3.2 Model optimisation
As mentioned above, a subset of background triggers and simulated WNB triggers are used to construct the background and signal GMM models respectively. While fitting mixture models to each of the two training sets, Gaussian parameters (mean, covariance and weight of each Gaussian) are optimised by the Expectation Maximisation (EM) algorithm Dempster et al. (1977), while the hyper-parameter of number of Gaussian components is specified by optimising detection efficiency with the validation data. During this validation step, models are produced over a range of number of Gaussian components. The log-likelihood values of all triggers (WNB and background) in validation data are calculated for each background and signal model over the range of the number of Gaussian components, allowing for the construction of the $T$ statistic as in eqn. 1. We define detection efficiency as the fraction of WNB validation triggers with $T$ value greater than the detection threshold at a given FAR. In general, the fraction of detected signals will depend on the waveform type and the injected amplitude. But for the detection efficiency calculation here, we use only the WNB triggers and consider all triggers above the detection threshold regardless of the amplitude. For model optimisation, we choose to calculate the detection efficiency at a FAR of 1 in 50 years for each combination of GMM models. The combination of models which give the optimal detection efficiency at the selected FAR threshold are labelled as the GMM models with the optimal number of Gaussian components. An example of the distribution of efficiency at FAR 1 in 50 years over the number of Gaussian components is seen for the O3a analysis in fig 1. GMMs were trained for number of Gaussian components in the range $[140,168]$ for background triggers and in the range $[120,188]$ for the WNB simulations, both varying in steps of 2. Only the region surrounding the optimal number of Gaussian components is shown in the figure. The colorbar represents the detection efficiency on validation data for each combination of models. We observe that the detection efficiency varies from $0.48$ to $0.52$ in this range of numbers of Gaussian components, with the maximum occurring at the combination of models with $148$ components for background and $172$ components for WNB simulations.
Previously in Lopez et al. (2022b), the optimal number of Gaussian components had been selected using the Bayesian Information Criterion (BIC). However, the BIC provided a measure of how well the Gaussians fit the training data, which did not directly relate to our models’ ability to distinguish between the background and signal populations. The new approach attempts to combine both measures through the optimisation of correctly classified data, and was found to reliably obtain better results.
@@ III.1 Comparisons with old GMM method
  • New method to optimise number of Gaussian components in the models: As demonstrated in Figure 1, this is now done through the optimisation of detection efficiency at an inverse False Alarm Rate (iFAR) threshold of 50 years over combinations of models with differing number of Gaussian components with validation data. Previously, the Bayesian Information Criterion (BIC) was utilized, however was not reliably favouring models with optimal classification performance.
@@
Figure 2: Percentage change in $h_{rss50}$ for the updated methodology, GMM+, compared to the old GMM methodology seen in Lopez et al. (2022b) for the generic set of ad-hoc waveform injections and CCSN injections at an iFAR threshold of 100 years. A negative change signifies lower $h_{rss50}$ and hence better sensitivities with the GMM+ methodology.
We directly compare the sensitivity of both methodologies to given waveforms by calculating the percentage change in $h_{rss50}$ at iFAR $\geq$ 100 years for GMM+ relative to the previous methodology, as seen in Figure 2. A negative percentage change indicates a lower $h_{rss50}$, and hence an improvement in sensitivity due to the ability to probe smaller amplitude signals for that given waveform. The results for ad-hoc simulation waveforms are shown in Fig 2(a). Sensitivities to the majority of Sine-Gaussian waveforms with $Q$=3, $Q$=9 are worsened, however this loss is somewhat expected due to the removal of bias from the models towards these specific injections. Sensitivities to Sine-Gaussian $Q=100$ and White Noise Burst waveforms remain comparable, while GMM+ gains sensitivity to the majority of Gaussian Pulse waveforms. The benefit of eliminating ad-hoc waveform bias from the model is further seen in Figure 2(b), where the consistent decrease in $h_{rss50}$ for CCSN injections demonstrates that our new GMM+ methodology improves the pipeline’s sensitivity to astrophysical source populations.
@@ V Conclusions
Considering the competitive sensitivities of the GMM and XGBoost post-productions to various astrophysical signals, it may in the future be desirable to run multiple pipelines on a dedicated search for GWs. The combination of multiple pipelines would require the application of a trials factor, however the true implications of this are not yet fully understood and not the target of this work.
@@ Acknowledgements.
The authors would like to thank Giovanni Prodi for useful comments. The authors would like to thank the authors of Szczepań czyk et al. (2023) for the production of cWB triggers used for the analysis in this paper. The authors would like to thank Gungwon Kang for interesting discussions. The authors acknowledge the computational resources which aided the completion of this project, provided by LIGO-Laboratory and support by the National Science Foundation (NSF) Grants No..PHY-0757058 and No.PHY-0823459. L.S acknowledges support from the Science and Technology Facilities Council [ST/V506692/1 2446745] and the PRIN project 202275HT58. SG acknowledges fellowship support from MHRD, Government of India. J. S. is supported in part by the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2022R1I1A207366012). G.V. acknowledge the support of the National Science Foundation under grant PHY-2207728. I.S.H. was supported by Science and Technology Facilities Council (STFC) grants ST/V001736/1 and ST/V005634/1. AP acknowledges the support from SPARC MoE grant SPARC/2019-2020/P2926/SL, Government of India.
@@ Appendix B CBC events detected by cWB-GMM in O3
Table 4 lists the CBC events detected by GMM+ with iFAR $\geq$ 1 year for both the LH and LHV networks. The iFAR estimates obtained with the other cWB pipelines are listed for the same events. The triggers are obtained from the entire coincident data used for the O3 cWB all-sky short search.